40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 605     Entry time: Mon Jun 30 15:56:22 2008
Author: Jenne 
Type: Update 
Category: Electronics 
Subject: Fixing the LO demod signal  
To make the alarm handler happy, at Rana and John's suggestion I replaced R14 of the MC's Demod board, changing it from 4.99 Ohms to 4.99 kOhms. This increased the gain of the LO portion of the demod board by a factor of 10. Sharon and I have remeasured the table of LO input to the demod board, and the output on the C1:IOO-MC_DEMOD_LO channel:

Input Amplitude to LO input on demod board [dBm]: | Value of channel C1:IOO-MC_DEMOD_LO
------------------------------------------------- | -----------------------------------
-10 | -0.00449867
-8 | 0.000384331
-6 | 0.0101503
-4 | 0.0296823
-2 | 0.0882783
0 | 0.2543
2 | 0.542397
4 | 0.962335
6 | 1.65572
8 | 2.34911
10 | 2.96925
ELOG V3.1.3-